OpenStudy (dls):

A boy stands 39.2m from a buliding and throws a ball which just passes through a window 19.6m above the ground.Calculate the velocity of projection of the ball..

OpenStudy (mayankdevnani):

kinematics......oh!!

OpenStudy (anonymous):

The "velocity of projection" part isn't really clear for me. What's that mean?

OpenStudy (anonymous):

\[v_x=vsin \theta\]\[v_y=vcos \theta\]

OpenStudy (anonymous):

\[s_x=vsin (\theta) t\] \[s_y=-\frac{1}{2}gt^2+vcos (\theta) t\]

OpenStudy (anonymous):

For\[s_x(39.2)\]Find the time, T.

OpenStudy (anonymous):

Then insert the value for T, and you have a system of 2 equations +2 unknowns. Substitute something to solve it.

OpenStudy (anonymous):

Use\[v=\frac{s_x}{\sin(\theta)t}\] and insert it into the second equation (or course, at T s_x=39.2 and s_y=19.6, so put those values in)